OpenStudy (anonymous):

Can someone help me with this problem please (km^1/2)^3 (k^3 m^5)^1/2 I need help with steps (km^1/2^)3 (k^3 m^5)^1/2 Power Rule (k^1/2)^3 (m^1/2)^3 (k^3)^1/2 (m^5)^1/2 Re applying the power rule (k^3/2) (m^3/2) (k^3/2) (m^5/2) Simplify k^3 m^4 Nth root

OpenStudy (anonymous):

can someone tell me if these are the correct steps. Please (km1/2)3 (k3 m5)1/2 Power Rule (k1/2)3 (m1/2)3 (k3)1/2 (m5)1/2 Re applying the power rule (k3/2) (m3/2) (k3/2) (m5/2) Simplify k3 m4 Nth root

OpenStudy (ranga):

(km^1/2)^3 (k^3 m^5)^1/2 = k^3 * m^(3/2) * k^(3/2) * m^(5/2) = k^(3 + 3/2) * m^(3/2 + 5/2) = k^(9/2) * m^(4)

OpenStudy (anonymous):

okay I see that my steps are off by a long shot.

OpenStudy (ranga):

In the 5th line in the original posting of the problem the mistake is k^1/2 on the left end. K is not raised to 1/2. It is raised to 3.

OpenStudy (anonymous):

Thank you for the help. I guess I need to study this more.

OpenStudy (ranga):

You got the "m" result correct. Just the k was a little off. You are welcome.